Studies of weather show that, historically, Jamaica is hit by a hurricane every 20 years. Calculate the following probabilities based on the historical record.

a. What is the probability that Jamaica will be hit by a hurricane in any given year? Show as a fraction and also show as a proportion with 2 decimal places. (To find the proportion in Excel, I recommend that you not use the Fraction data type. For this problem, divide 1 by the number of years)

b. What is the probability that Jamaica will NOT be hit by a hurricane in any given year? Show as a fraction and also show as a proportion with 2 decimal places.
Answer #1

(a) answer is =1/20=0.05

(b) here we can use poisson distribution with mean=\lambda=1 ( as the probability of this event is very less)

for poisson distribution P(X=x)=exp(-\lambda)\lambda^x/x!

for P(X=0)=exp(-1)10/0!=exp(-1)=0.3679